# Private Linear Regression with Differential Privacy and PAC Privacy

### v1

@@ 3.1 Differential Privacy
[…]
@@ 3.2 PAC Privacy
To align with the existing privacy notions like differential privacy, we focus on the special case of PAC privacy: membership privacy [32]. Specifically, let $\bm{1}_{i}$ represent an indicator which equals $1$ if the $i$-th datapoint $u_{i}$ from the entire data pool $\mathsf{U}=\{u_{1},u_{2},\cdots,u_{n}\}$ is included in the sampled set $X$ otherwise 00. The adversary aims to recover $\bm{1}_{i}$ by some estimate $\hat{\bm{1}}_{i}$. As a corollary from [30], under a Poisson sampling $q$, the mutual information $\mathsf{MI}\big{(}\bm{1}_{i},\mathcal{M}(X)\big{)}$ can be upper bounded as
@@
Here, $X_{-i}$ represents a random set produced by conducting a Poisson sampling of rate $q$ over $\mathsf{U}/u_{i}$. In particular, when we consider adding (multivariate) Gaussian noise to some deterministic processing function $\mathcal{F}(\cdot)$, i.e., $\mathcal{M}(X)=\mathcal{F}(X)+e$ for $e\sim\mathcal{N}(0,\Sigma)$, then
@@ 4.3 Linear Regression with PAC Privacy
         if mean of $v[i]-$ previous mean $<\eta$ then
@@
      Compute noise For $\text{For $i\in[d]$, }e[i]\leftarrow\frac{\sqrt{v[i]}\sum_{j=1}^{d}\sqrt{v[j]}}{4\cdot\text{MI}}$ Project to original space $d\times d$ diagonal matrix $\Sigma^{\prime}$: $\Sigma^{\prime}[i][i]\leftarrow e[i]$ for $i\in[d]$ $\Sigma=V^{T}\cdot\Sigma^{\prime}$ $e[i]\leftarrow\Sigma[i][i]$ for $i\in[d]$ $e_{max}[i]=\max(e[i],e_{max}[i])$ for $i\in[d]$
@@ 5.1 Experimental Setups
Lenses $24$ $4$ Regression
Concrete $103$ $7$ Regression

### v2

@@ 3.1 Differential Privacy
Define a randomized mechanism $\mathcal{M}\penalty 10000\mskip 2.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 6.0mu plus 1.0mu\mathcal{X}\to\mathcal{Y}$. Then, for two datasets $D,D^{\prime}\subseteq\mathcal{X}$ that differ by the presence or absence of a single record, $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private if for any $y\subseteq\mathcal{Y}$
@@ 3.2 PAC Privacy
To align with the existing privacy notions like differential privacy, we focus on the special case of PAC privacy: membership privacy [32]. Specifically, let $\bm{1}_{i}$ represent an indicator which equals $1$ if the $i$-th datapoint $u_{i}$ from the entire data pool $\mathsf{U}=\{u_{1},u_{2},\cdots,u_{n}\}$ is included in the sampled set $X$ otherwise 0. The adversary aims to recover $\bm{1}_{i}$ by some estimate $\hat{\bm{1}}_{i}$. As a corollary from [30], under a Poisson sampling $q$, the mutual information $\mathsf{MI}\big(\bm{1}_{i},\mathcal{M}(X)\big)$ can be upper bounded as
@@
Here, $X_{-i}$ represents a random set produced by conducting a Poisson sampling of rate $q$ over $\mathsf{U}/\penalty 50u_{i}$. In particular, when we consider adding (multivariate) Gaussian noise to some deterministic processing function $\mathcal{F}(\cdot)$, i.e., $\mathcal{M}(X)=\mathcal{F}(X)+e$ for $e\sim\mathcal{N}(0,\Sigma)$, then
@@ 2 Anisotropic noise estimation for PAC membership privacy
   if mean of $v[i]$- previous mean $\eta$ then
@@
   Compute noise $\text{For $i\in[d]$, }e[i]\leftarrow\frac{\sqrt{v[i]}\sum_{j=1}^{d}\sqrt{v[j]}}{4\cdot\text{MI}}$ Project to original space $d\times d$ diagonal matrix $\Sigma^{\prime}$: $\Sigma^{\prime}[i][i]\leftarrow e[i]$ for $i\in[d]$ $\Sigma=V^{T}\cdot\Sigma^{\prime}$ $e[i]\leftarrow\Sigma[i][i]$ for $i\in[d]$ $e_{max}[i]=\max(e[i],e_{max}[i])$ for $i\in[d]$
@@ 5.1 Experimental Setups
Lenses $24$ 44 Regression
Concrete $103$ 77 Regression